# DDN-SLAM: Real-time Dense Dynamic Neural Implicit SLAM with Joint Semantic Encoding

### 03 Jan 2024 v1

@@ 5 Conclusion
We propose DDN-SLAM, a neural implicit semantic SLAM system that enables stable tracking in complex occlusion and challenging environments. It achieves dense indoor scene reconstruction through 2D-3D scene understanding. Our system is more aligned with real-world requirements and can flexibly support various inputs while demonstrating superior performance. It accurately captures geometric features and texture details, while enabling long-term updates and global optimization. Our experiments demonstrate that our method effectively avoids interference from severe occlusion on mapping and tracking, and achieves state-of-the-art performance on various datasets. Limitations. The application of our system on mobile platforms such as robot platforms has promising potential. Furthermore, more reasonable depth estimation and detail optimization with spatio-temporal consistency may lead to better mapping results.

### 09 Mar 2024 v2

@@ 5 Conclusion
We propose DDN-SLAM, a neural implicit semantic SLAM system that enables stable tracking in complex occlusion and challenging environments. Through 2D-3D scene understanding, it achieves dense indoor scene reconstruction. Our system is more aligned with real-world requirements, capable of flexibly supporting various inputs while demonstrating superior performance. It addresses occlusion from complex moving objects, retains and reconstructs potential moving objects in the scene, all while balancing computational memory and running speed. Our experiments demonstrate that our method effectively avoids interference from severe occlusion on mapping and tracking, and achieves state-of-the-art performance on various datasets.